OpenStudy (thomasfire55):

Rewrite the quadratic function in vertex form. Then, determine the maximum or minimum and the axis of symmetry. y = -3x^2 + 18x - 2

OpenStudy (photon336):

Well, to start this is what you would need. \[y = a(x-h)^{2}+k\] \[Axis~of~symmetry = \frac{ -b }{ 2a }\]

OpenStudy (thomasfire55):

whats a,h,k?

OpenStudy (thomasfire55):

like what numbers are they in this

OpenStudy (triciaal):